OpenStudy (anonymous):

-3 ≤ 7c + 4 < 18 Show steps on how you got the answer, thanks!! edited question bc of typo D':

OpenStudy (texaschic101):

there is 2 ways to do this...you can break this apart into 2 inequalities.. -3 <= 7c + 4 and 7c + 4 > 8.....and solve them both...or you can get c by itself...I would have to show you how to do that. So which way do you want to proceed ?

OpenStudy (anonymous):

Could you show me how to do both?

OpenStudy (texaschic101):

-3 <= 7c + 4 7c + 4 > 8 -3 - 4 <= 7c 7c > 8 - 4 -7 <= 7c 7c > 4 -1 <= c c > 4/7 and when put together, you get : -1 <= c > 4/7 Or this way... -3 <= 7c + 4 > 8 --- subtract 4 from all 3 -4 -3 <= 7c + 4 - 4 > 8 - 4 -- simplify -7 <= 7c > 4 -- divide all sides by 7 -7/7 <= 7c/7 > 4/7 -- simplify -1 <= c > 4/7

OpenStudy (anonymous):

thank you!!!!!

OpenStudy (texaschic101):

sure thing :)

OpenStudy (freckles):

the inequality makes no sense -3>8 is not true

OpenStudy (freckles):

-3 ≤ 7c + 4 > 8 was this meant to be : -3 ≤ 7c+4 < 8 ?

OpenStudy (anonymous):

Oh!! Sorry I just realized the typo!! The problem's actually -3 ≤ 7c + 4 < 18. D':